Parth (parthkohli):

Show that if \(1 = as + bt\), \(\gcd(a,b) = 1\)

Parth (parthkohli):

So actually is it just the Bezout's Lemma?

Parth (parthkohli):

\[\gcd(a,b) = as + bt \]

Parth (parthkohli):

\[1 = \gcd(a,b) = as + bt\]So,\[1 = \gcd(a,b)\]That's the only work?

OpenStudy (unklerhaukus):

greatest common divisor?

Parth (parthkohli):

yes

OpenStudy (anonymous):

its a theorem of number theory right?

Parth (parthkohli):

Yes, elementary number theory :)

OpenStudy (anonymous):

ya..........its a result that when and b are relatively prime .....then there exists integers s and t such that 1=as +bt

OpenStudy (anonymous):

gcd(a,b) = 1, then gcd(a^2,b^2)=1. Since both = 1 we can write this as gcd(a,b) = gcd(a^2,b^2) Then a = a^2 and b = b ^2 There are only 2 real numbers that when squared equals itself, 0 and 1. Since we are assuming that both equations in fact = 1, then a and b both equal one, If either gcd (a,b) or gcd(a^2,b^2)= 0 the only other choice, then gcd (a,b) or gcd(a^2,b^2) would not equal 1.

Parth (parthkohli):

What? o.O

Parth (parthkohli):

I asked if my statement was enough. lol

OpenStudy (experimentx):

special case of Bezuot's identity.

Parth (parthkohli):

Exactly -- can you see my solution?

OpenStudy (experimentx):

looks like both works either way!! that should be right!!

Parth (parthkohli):

Okay :)

OpenStudy (experimentx):

although proving this lemma is fairly complicated.